### v1

# LayerPano3D: Layered 3D Panorama for Hyper-Immersive Scene Generation

Abstract
3D immersive scene generation is a challenging yet critical task in computer vision and graphics. A desired virtual 3D scene should 1) exhibit omnidirectional view consistency, and 2) allow for free exploration in complex scene hierarchies. Existing methods either rely on successive scene expansion via inpainting or employ panorama representation to represent large FOV scene environments. However, the generated scene suffers from semantic drift during expansion and is unable to handle occlusion among scene hierarchies. To tackle these challenges, we introduce LayerPano3D, a novel framework for full-view, explorable panoramic 3D scene generation from a single text prompt. Our key insight is to decompose a reference 2D panorama into multiple layers at different depth levels, where each layer reveals the unseen space from the reference views via diffusion prior. LayerPano3D comprises multiple dedicated designs: 1) we introduce a novel text-guided anchor view synthesis pipeline for high-quality, consistent panorama generation. 2) We pioneer the Layered 3D Panorama as underlying representation to manage complex scene hierarchies and lift it into 3D Gaussians to splat detailed 360-degree omnidirectional scenes with unconstrained viewing paths. Extensive experiments demonstrate that our framework generates state-of-the-art 3D panoramic scene in both full view consistency and immersive exploratory experience. We believe that LayerPano3D holds promise for advancing 3D panoramic scene creation with numerous applications.
Index Terms:
3D Scene Generation, Panorama Generation, Diffusion Models, Neural Rendering
Figure 1: Overview of LayerPano3D. Guided by simple text prompts, LayerPano3D leverages multi-layered 3D panorama to create hyper-immersive panoramic scene with $360^{\circ}\times 180^{\circ}$ coverage, enabling free 3D exploration among complex scene hierarchies.
1 Introduction
The development of spatial computing, including virtual and mixed reality systems, greatly enhances user engagement across various applications, and drives demand for explorable, high-quality 3D environments. We contend that a desired virtual 3D scene should 1) exhibit high-quality and consistency in appearance and geometry across the full $360^{\circ}\times 180^{\circ}$ view; 2) allow for free exploration among complex scene hierarchies with clear parallax. In recent years, many approaches in 3D scene generation [1, 2, 3] were proposed to address these needs.
One branch of works [4, 5, 6, 7, 8, 9] seeks to create extensive scenes by leveraging a “navigate-and-imagine” strategy, which successively applies novel-view rendering and outpaints unseen areas to expand the scene. However, this type of approaches suffer from the semantic drift issue: long sequential scene expansion easily produces incoherent results as the out-paint artifacts accumulate through iterations, hampering the global consistency and harmony of the generated scene.
Another branch of methods [10, 11, 12, 13, 14, 15] employs Equirectangular Panorama to represent $360^{\circ}$, large field of view (FOV) environments in 2D. However, the absence of large-scale panoramic datasets hinders the capability of panorama generation systems, resulting in low-resolution images with simple structures and sparse assets. Moreover, 2D panorama [10, 11, 13] does not allow for free scene exploration. Even when lifted to a panoramic scene [16], the simple spherical structure fails to provide complex scene hierarchies with clear parallax, leading to occluded spaces that cause blurry renderings, ambiguity, and gaps in the generated 3D panorama. Some methods [17] typically use inpainting-based disocclusion strategy to fill in the unseen spaces, but they require specific, predefined rendering paths tailored for each scene, limiting the potential for free exploration.
To this end, we present LayerPano3D, a novel framework that leverages Multi-Layered 3D Panorama for full-view consistent and free exploratory scene generation from text prompts. The main idea is to create a Layered 3D Panorama by first generating a reference panorama and treating it as a multi-layered composition, where each layer depicts scene content at a specific depth level. In this regard, it allows us to create complex scene hierarchies by placing occluded assets in different depth layers at full appearance.
Our contributions are two-fold. First, to generate high-quality and coherent $360^{\circ}\times 180^{\circ}$ panoramas, we propose a novel text-guided anchor view synthesis pipeline. By finetuning a T2I model [18] to generate 4 orthogonal perspective views as anchors, we prevent semantic drifts during panorama generation, while ensuring a consistent horizon level across all views. Furthermore, the anchor views enrich the panorama by incorporating complex structures and detailed features derived from large-scale, pre-trained perspective image generators. Second, we introduce the Layered 3D Panorama representation as a general solution to handle occlusion for different types of scenes with complex scene hierarchies, and lift it to 3D Gaussians [19] to enable free 3D exploration. By leveraging pre-trained panoptic segmentation prior and K-Means clustering, we streamline an automatic layer construction pipeline to decompose the reference panorama into different depth layers. The unseen space at each layer is synthesized with a finetuned panorama inpainter [11].
Extensive experiments demonstrate the effectiveness of LayerPano3D in generating hyper-immersive layered panoramic scene from a single text prompt. LayerPano3D surpasses state-of-the-art methods in creating coherent, plausible, text-aligned 2D panorama and full-view consistent, explorable 3D panoramic environments. Furthermore, our framework streamlines an automatic pipeline without any scene-specific navigation paths, providing more user-friendly interface for non-experts. We believe that LayerPano3D effectively enhances the accessibility of full-view, explorable AIGC 3D environments for real-world applications.
2 Related Works
In recent years, advancements in AI, particularly deep learning, along with increased computational power and large datasets, have driven rapid development in 2D and 3D generation  [20]. These technologies have revolutionized digital content creation in fields like entertainment, design, and virtual reality.
2.1 3D Representation
Early neural rendering methods [21, 22, 23, 24] typically incorporate a multi-layer perceptron (MLP) to model 3D scenes as continuous functions. NeRF [25], being the representative MLP-based method, achieves state-of-the-art novel view synthesis performance due to the volume rendering and inductive bias nature of MLPs. However, as the MLP needs to evaluate on large amount of points along every camera ray, the rendering process becomes extremely exhaustive. Subsequent methods [26, 27, 28, 29, 19] improve on faster training and rendering by employing sparse, advanced scene representation. Among these, 3D Gaussian Splatting (3DGS) [19] has recently emerged as a rising star with transformative real-time rendering and high-definition resolutions. It explicitly parameterizes the 3D scene with 3D Gaussians, where each Gaussian is optimized to represent a volume and projected to 2D for rasterization. The discrete nature of 3D Gaussians enables flexible scene editing and composition for 3DGS. Therefore, our approach employs 3DGS as scene representation to facilitate multi-layer scene creation.
2.2 3D Scene Generation
Due to the recent success of diffusion models, 3D scene generation has also achieved some development. Scenescape [7] and DiffDreamer [30], for example, explore perpetual view generation through the incremental construction of 3D scenes. One major branch of work employ step-by-step inpainting from pre-defined trajectories. Text2Room [6] creates room-scale 3D scenes based on text prompt, utilizing textured 3D meshes for scene representation. Similarly, LucidDreamer [4] and WonderJourney [5] can generate domain-free 3D Gaussian splatting scenes from iterative inpainting. However, this line of work often suffer from the semantic drift issue, resulting in unrealistic scene from artifact accumulation and inconsistent semantics. While some other approaches [31, 3, 32] endeavor to integrate objects with environments, they yield relatively low quality of comprehensive scene generation. Recently, our concurrent works, DreamScene360 [16] and HoloDreamer [17] also employ panorama as prior to construct panoramic scenes. However, they only achieve the $360^{\circ}\times 180^{\circ}$ field of view at a fixed viewpoint based on a single panorama of low-quality and simple structure, and do not support free roaming within the scene. In contrast, our framework leverages Multi-Layered 3D Panorama representation to construct high-quality, fully enclosed scenes that enable unconstrained navigation paths in 3D scene.
2.3 Panorama Generation
Panorama generation methods are often based on GANs or diffusion models. Early in this field, with the different forms of deep generative neural networks, GAN-based panorama generation methods explore many paths to improve quality and diversity. Among them, Text2Light [15] focuses on HDR panoramic images by employing a text-conditioned global sampler alongside a structure-aware local sampler. However, training GANs is challenging and they encounter the issue of mode collapse. Recently, some studies have utilized diffusion models to generate panoramas. MVDiffusion [10] generates eight perspective views with multi-branch UNet but the resulting closed-loop panorama only captures the $360^{\circ}\times 90^{\circ}$ FOV. The image generated from MultiDiffusion [33] and Syncdiffusion [34] is more like a long-range image with wide horizontal angle as they do not integrate camera projection models. PanoDiff [35] can generate $360^{\circ}$ panorama from one or more unregistered Narrow Field-of-View (NFoV) images with pose estimation and controlling partial FOV LDM, while the quality and diversity of results are limited by the scarcity of panoramic image training data like most other methods [14, 36, 37]. In contrast, our model can generate Multi-Layered 3D Panorama for immersive, high-quality, and coherent scene generation from text prompts.
3 Method
Figure 2: Pipeline Overview of LayerPano3D. Our framework consists of three stages, namely reference panorama generation, multi-layer panorama construction and panoramic 3D scene optimization.  LayerPano3D streamlines an automatic generation pipeline without any manual efforts to design scene-specific navigation paths for expansion or completion.
The goal of our work is to create a panoramic scene guided by text prompts, that encompasses a complete $360^{\circ}\times 180^{\circ}$ field of view from various viewpoints within an extensive range in the scene, while allowing for unconstrained trajectory for immersive exploration. LayerPano3D consists three stages. In Stage I (Sec. 3.1), we propose a text-guided anchor view synthesis pipeline paired with panoramic outpainting to generate high-quality, consistent panorama as reference. In Stage II (Sec. 3.2), with the reference panorama, we construct our Layered 3D Panorama representation by iterative layer decomposition, completion and alignment process. In Stage III (Sec. 3.3), the Layered 3D Panorama is lifted to 3D Gaussians in a cascaded manner to enable free 3D exploration.
3.1 Reference Panorama Generation
Text-Guided Anchor View Synthesis. We start by generating four orthogonal anchor views to establish the fundamental geometric structure and visual appearance. This is achieved by fine-tuning a pre-trained perspective T2I diffusion model.
For data preparation, we construct a high-quality RGB panorama dataset consisting of 911 outdoor images sourced from the web and 9684 images from Matterport3D [38]. For each panorama, we use llava-v1.5-7b [39] to generate captions as text prompts and project four orthogonal views without overlap at a fixed FOV ($60^{\circ}$), elevation ($0^{\circ}$) and four azimuths ($0^{\circ},90^{\circ},180^{\circ},270^{\circ}$). This setting can effectively avoid perspective distortion and fisheye effect, and allow further refinement of consistency in the subsequent synthesis stages. Additionally, we arrange these four views into a $2\times 2$ grid image in a fixed order for the ease of training and draw a white cross-line with a certain pixel size between the grids as a split.
Compared with object-level multi-view generation, it is harder for scene-level generation to diffuse more views with high quality and consistency in a single pass. Multi-view cross-attention-based methods [10, 11] learn high-level semantics at the expense of overfitting to monotonous patterns due to the scarcity of panorama data. Other approaches [1] require huge computation and time in both training and inference stages and generate relatively simple scenes with limited assets. Therefore, inspired by [40], we directly fine-tune Stable-Diffusion-XL [18] to efficiently generate high-quality anchor views as image grids for subsequent panorama generation, free of complex regularization techniques. This text-guided anchor view synthesis pipeline can scale positively with the potency of the underlying T2I model and unlock the capability of the base model without additional knowledge through lightweight fine-tuning. Furthermore, unlike the clean background in the object-level generation, scene backgrounds are often diverse and complex. Thus, instead of simply initializing each sub-grid with a 2D Gaussian blob as is commonly done in object-based multi-view approaches, we initialize the image grids with random colored noise and draw a white cross-line between the grids during the inference process, mimicking the training data.
Equirectangular Panorama Synthesis. Equirectangular Projection (ERP) is a method that maps a 3D sphere onto a 2D map. In ERP, the lines of longitude are mapped to vertical lines with constant spacing, and the lines of latitude are mapped to horizontal lines with constant spacing. This mapping relationship is neither area-preserving nor angle-preserving and the mapping of latitude leads to strong distortions in the polar regions. To obtain a detailed, consistent and plausible panorama, we first generate a $360^{\circ}\times 90^{\circ}$ FOV panorama to obtain the principle structure and content, then successively extend it to $360^{\circ}\times 180^{\circ}$ FOV panorama.
With the text-guided anchor view synthesis, we have the anchor perspective images $[I_{anc}^{i}]_{i=1}^{4}\in\mathbb{R}^{h\times w\times 3}$ of a scene captured by four orthogonal camera in the horizontal plane. Based on the mapping rule and $60^{\circ}$ FOV, these anchor views can be directly projected to form the incomplete $360^{\circ}\times 60^{\circ}$ FOV panorama without overlap. Here, we assume that the elevation angle of the anchors are $\phi_{anc}=0^{\circ}$ with azimuth angles $\{\theta_{anc}^{i}\}=\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$. Then, we set a viewpoint sequence with $\{(\theta_{j},\phi_{j})\}_{j=1}^{n}$ that allows for a 360-degree rotation around the scene in $90^{\circ}$ FOV to project the perspective views $\hat{I_{j}}$ and masks $M_{j}$:
$[\hat{I_{j}},M_{j}]=\mathcal{ERP}^{-1}(Pano,[\theta_{j},\phi_{j},FOV]),$ (1)
where the $\mathcal{ERP}^{-1}$ is the inverse ERP from panorama to perspective views in 3D. We further apply a text-guided inpainter $\mathcal{F}_{inpaint}$ [41] to synthesize the missing pixels in the masked regions $M_{j}$ of the incomplete views $\hat{I_{j}}$. Note that, $\hat{I_{j}}$ is additionally pre-processed with LaMa [42] before inpainting to clean up the masked pixel distribution with contextual knowledge:
$I_{j}=\mathcal{F}_{inpaint}(LaMa(\hat{I_{j}},M_{j})).$ (2)
Through iterative projection and inpainting, the $Pano_{360\times 60}$ is extended to $Pano_{360\times 90}$.
The extension from $Pano_{360\times 90}$ to $Pano_{360\times 180}$ is more difficult, as the polar regions exhibit most distorted patterns that cannot simply be synthesized through perspective image inpainting. Therefore, we fine-tune [43] on panoramic image-caption pairs for polar outpainting. Another challenge is to enforce continuity between the leftmost and rightmost sides of the synthesized image during outpainting, as it is a crucial characteristic [44] of 360 panoramas. Following  [14, 33], we adopt a circular blending strategy in both denoising and VAE decoding stage, as illustrated in Figure 3. A circular closed loop is created by extending the rightmost side of the panorama with the pixels in the leftmost parts. Per-pixel adaptive weights $W_{k}$ are utilized to blend the denoised patches mapped from extensions and from the leftmost parts to generate the leftmost panorama. Similarly, in the VAE tiled decoding stage, to avoid tiling artifacts, we ensure that each tile overlaps and blends together to form a smooth output due to each tile using a different decoder. Therefore, each pixel of the panorama is computed as a weighted average of all its mapped patch updates $Z_{k,t}$ at each denoising step.
$\Psi(\mathcal{P}_{t})=\sum_{k}\frac{\mathcal{F}_{map,k}^{-1}(W_{k})}{\sum_{m}\mathcal{F}_{map,m}^{-1}(W_{m})}\odot\mathcal{F}_{map,k}^{-1}(\Phi(Z_{k,t})),$ (3)
where $\mathcal{F}_{map,k}^{-1}$ is image space mapping from patch $Z_{k}$ to its corresponding latent in $\mathcal{P}$; $\Psi$ and $\Phi$ denote the diffusion process at panorama-level and patch-level. Finally, we can get a high-quality result as reference for the subsequent layered panorama construction.
Figure 3: The circular blending Strategy. During the inference process, to synthesize a $360^{\circ}$ seamless panorama, adaptive weights are utilized to blend the leftmost part of the latent with the rightmost part in denoising and VAE decoding stage.
3.2 Multi-Layer Panorama Construction
We introduce the Layered 3D Panorama representation based on the following assumption: “an enclosed 3D scene contains a background and various assets positioned in front of it”. In this regard, using Layered 3D Panorama for 3D scene generation is a general approach to handle occlusion for various types of scene. To create a complete scene from the reference panorama, we propose to decompose it into $N+1$ layers along the depth dimension. As shown in Figure 2, these layers, arranged from farthest to nearest, represent both the scene background (layer 0) and the layouts situated behind the observation point.
Layer Decomposition. As shown in Figure 2, the reference panorama is decomposed by first identifying the scene assets and then cluster these assets in different layers according to depth. First, we employ an off-the-shelf panoptic segmentation model [45] pretrained on ADE20K [46] to automatically find all scene assets visible in the reference panorama. A good layer decomposition requires that the layer assets share a similar depth level within layers and are distant from assets in other layers. In this sense, we assign each asset a depth value and apply K-Means to cluster these masks into different groups. Given the reference panorama depth map, the depth value for each asset mask is determined by calculating the 75th percentile of the depth values within the masked region. According to the depth values, the assets are clustered into $N$ groups from layer 00 to $N-1$ and are merged into layer masks to guide the subsequent layer completion.
Layer Completion. With the layer mask, we focus on completing the unseen content caused by asset occlusion. In order to synthesize background pixels instead of creating new elements, we finetune PanFusion [11] with the ERP-perspective aware cross-attention as the panoramic background inpainter. Specifically, at each layer, our model takes the layer mask $M_{l}$, the reference panorama, and the “empty scene, nothing” [47] prompt as input, and output coherent content at the masked area. The inpainted panorama at layer $l$ is denoted $P_{l}$ and is used as supervision to the subsequent panoramic 3D Gaussian scene optimization. Note that, we additionally apply SAM [48] to extend the layer mask, based on the inpainted panorama from the previous layer, to eliminate unwanted new generations from inpainting.
Moreover, to enable free rendering in 3D, where observers can examine scenes from varying distances, the unprocessed textures of distant assets may appear blurred as the observer approaches. Therefore, distant layers require higher resolution to preserve texture details at different viewpoints. To address this, Super Resolution (SR) module [49] is employed to enhance the resolution of the layered panorama from layer 00 (background layer) to layer $N$ (reference panorama), achieving a 2×2$\times$2 × upscale in resolution. SR processing significantly improves the texture quality of distant objects, maintaining their visual clarity and texture details even when observed from a closer perspective.
Layer Alignment. Given the Layered 3D RGB Panorama $[P_{l}]_{l=0}^{N}$, we perform the depth prediction and alignment to ensure consistency in a shared space. To begin with, we apply the 360MonoDepth [50], to first estimate the layer $N$ (reference panorama) as the reference depth $P_{depth}^{N}$. The predicted depth $P_{depth}$ inherently contains extremely small values from normalization, causing peak-like artifacts under the reference camera position. Consequently, we post-process the directly output disparity map $\hat{P}_{disp}$ with the following transformation to mitigate the geometry distortion:
$D_{\text{bias}}$ $=\left(\max(\hat{P}_{\text{disp}})-\min(\hat{P}_{\text{disp}})\right)\ /\ 4,$ (4)
$\hat{P}_{\text{depth}}$ $=\max(\hat{P}_{\text{disp}})-\hat{P}_{\text{disp}}+D_{\text{bias}},$
$P_{\text{depth}}$ $=\hat{P}_{\text{depth}}\ /\ \max(\hat{P}_{\text{depth}}).$
To align the layer depth in 3D space, we find it infeasible to simply compute a global shift and scale as in  [4, 6] due to the nonlinear nature of ERP. Therefore, we leverage depth inpainting model $\mathcal{F}_{depth}$ from [51] to directly restore depth values based on the inpainted RGB pixels. $\mathcal{F}_{depth}$ harnesses strong generalizability from large-scale diffusion prior and synthesizes inpainted depth values at an aligned scale with the base depth. We start from reference panoramic depth $P_{depth}^{N}$ to implement step-by-step restoration from layer $N-1$ to layer 00:
$P_{depth}^{l}=\mathcal{F}_{depth}(P_{l},M_{l}\odot P_{depth}^{l+1}),$ (5)
where, in layer $l$, the inpainted panorama $P_{l}$ and masked depth map $M_{l}\odot P_{depth}^{l+1}$ are provided as inputs to $\mathcal{F}_{depth}$ for restoration.
3.3 Panoramic 3D Gaussian Scene Optimization
Figure 4: Illustration of the Gaussian Selector. Given the new asset point cloud, the Gaussian Selector identifies the active Gaussians for next layer’s optimization. (a) Optimized Gaussians at layer $l$ partially blocks new asset point cloud in layer $l+1$. (b) By computing the viewing directions and the distance to camera, Gaussian Selector identifies the blocking Gaussians. (c) The marked Gaussians are labeled as active, and merged with the new points for optimization. (d) Showed a special case where the optimized Gaussian center is behind the new points, but also identifies as an active Gaussian because its volume scale overlaps with the points (case 1).
3D Scene Initialization. To enable free 3D exploration, we lift the Layered 3D Panorama to 3D Gaussians [19], where the Gaussians are initialized from the layered 3D panoramic point clouds. Considering the intrinsic spherical structure of panorama, we can easily transform an equirectangular image $P\in\mathbb{R}^{H\times W\times 3}$ into 3D point cloud $S(\theta,\phi,P_{depth})$. Each pixel $(u,v)$ is represented as a 3D point and the angles $\theta,\phi$ are computed as $\theta=(2u/W-1)\pi$, $\phi=(2v/H-1)\pi/2$.
Then, the corresponding 3D coordinates $(X,Y,Z)$ from the depth value $P_{depth}(\theta_{u},\phi_{v})$ are derived as follows:
$X$ $=P_{depth}(\theta_{u},\phi_{v})\cos{\phi_{v}}\cos{\theta_{u}},$ (6)
$Y$ $=P_{depth}(\theta_{u},\phi_{v})\sin{\phi_{v}},$
$Z$ $=P_{depth}(\theta_{u},\phi_{v})\cos{\phi_{v}}\sin{\theta_{u}}.$
Based on this transformation, we can extract the point cloud for each layer panorama to initialize 3D Gaussians.
Drastic depth changes at layout edges introduce noisy stretched outliers that would turn into artifacts during scene refinement. Therefore, we propose an outlier removal module that specifically targets stretched point removal using heuristic point cloud filtering strategies. As stretched points are usually sparsely distributed in space, we design the point filtering strategy based on its distance from the neighbors. First, we filter out all points with the minimum distance to neighbors over threshold $\beta_{1}$. Then, we eliminate points with very few neighbors. The idea is to calculate the number of neighbors of each point within a given radius and drop the points where their number of neighbors is below threshold $\beta_{2}$. To speed up the calculation, we map the points into 3D grids, then remove all points within grids that have less than $\beta_{2}$ number of neighbors.
3D Scene Refinement. During scene refinement, we devise two types of Gaussian training schemes for varying scene content: the base Gaussian for reconstructing the scene background and the layer Gaussian for optimizing scene layouts. Additionally, a Gaussian selector module is introduced between layer Gaussians to facilitate scene composition.
In scene refinement, the base Gaussian model is initialized on a whole of the background point cloud, and the layer Gaussian model initiates on and optimizes the foreground assets. In practice, we project the layer mask $\mathcal{M}_{l}$ onto point clouds and use the masked points to initiate Gaussians. The optimized Gaussians from previous layers are frozen to avoid unwanted modification. In this way, the scene background is optimized once in the base Gaussian to reduce unnecessary computation and conflicts of Gaussians in subsequent layers.
We observe that the quality of the optimized scene is easily hampered by unaligned layers, and sometimes $\mathcal{F}_{depth}$ does fail to produce perfectly aligned layer depths. Gaussians at layer $l$ could span into unwanted depth levels and block assets in the subsequent layer, as illustrated in Figure 4(a). To handle this issue, we introduce the Gaussian selector module to detect these conflicted Gaussians, re-activate them from frozen, and optimize them away from the blockage. First, the selector computes the distance vector from the camera center $\mathbf{o}=(0,0,0)$ to each new point $\mathbf{p}$, as in Figure 4(b). The absolute distance from asset points $\mathbf{p}$ and scene Gaussians $\mathbf{g}$ to the camera is denoted as $d_{\mathbf{p}}$ and $d_{\mathbf{g}}$ respectively:
$d_{\mathbf{p}}=||\mathbf{p}-\mathbf{o}||_{2},\quad d_{\mathbf{g}}=||\mathbf{g}-\mathbf{o}||_{2}.$ (7)
By examining all Gaussians that on the same ray with asset points but at a closer distance: $\mathbf{p}\ /\ d_{\mathbf{p}}=\mathbf{g}\ /\ d_{\mathbf{g}},\quad d_{\mathbf{g}}<d_{\mathbf{p}},$ we mark them as active (Figure 4(c)). We also observe that despite the Gaussian center being behind the asset points, if the scaling of the Gaussian covers the point, then the Gaussian is still affecting the overall rendering quality. Hence, we update the gaussian distance as
$d^{\prime}_{\mathbf{g}}=d^{\prime}_{\mathbf{g}}-\max\{\mathbf{s}(\mathbf{g})\},$ (8)
where $\mathbf{s}$ is the scaling of the Gaussian $\mathbf{g}$. If $d_{\mathbf{g_{1}}}>d_{\mathbf{p_{1}}}$ and $d^{\prime}_{\mathbf{g_{1}}}<d_{\mathbf{p_{1}}}$ as in Figure 4(d), the referred Gaussian is also marked active. For efficient memory storage and fast look-up, we hash the distance vectors into a 3D grid. The mapping function from vector coordinates to grid indices writes: $f(\mathbf{p})=\text{ceil}(\beta_{3}\log(\textbf{p}+1))$.
Figure 5: Qualitative comparisons in Panorama Generation. We compare 2D panorama generation methods across three dimensions: Creativity (red), Resolution (green), and Spherical Structure (blue). LayerPano3D produces high-resolution, creative, and consistent panoramas while accommodating multi-style generation requirements.
4 Experiments
4.1 Implementation Details
In the anchor view synthesis stage, we train the model starting from SDXL-base-1.0 [18] with a batch size of 16 and learning rate of $10^{-5}$ for only 10K iterations on the 11k data pairs. The training is done using 4 NVIDIA A100 GPUs for 7 hours. For inference, we initialize four $512\times 512$ grids with random colored noise, subsequently superimposing a white cross-line with a pixel width of 10 to ensure that the final inference result can be accurately segmented into four grids of size $507\times 507$. In the reference panorama synthesis stage, we leverage the Stable Diffusion Inpainting pipeline [41] from huggingface [52] as the inpainter and repurposed Diffusion360 [43] as the outpainter to synthesize $360^{\circ}\times 180^{\circ}$ FOV panorama.
In the layered panorama construction stage, we employ OneFormer [45] to obtain the panoptic segmentation map for the reference panorama. Background categories are manually determined (i.e. sky, floor, ceiling, etc.) to filter out background components in asset masks. Generally, we cluster all asset masks into $N=3$ layers via KNN and merge all masks within each layer to form a unified layer mask. With the obtained layer mask, we combine our fine-tuned panorama inpainting model with LaMa [42] to achieve multi-layer completion and apply 360MonoDepth [50] to predict the reference panorama depth.
In the 3D panoramic scene optimization stage, we lift the panorama RGBD into 3D point clouds. For scene initialization, we set $\beta_{1}$ to $0.0001$ and $\beta_{2}$ to $4$ based on empirical practice. These point clouds are used to initialize the base Gaussian model and the layer Gaussian model. During the scene refinement stage, we optimize the base Gaussian model for 3,000 iterations, then the layer Gaussians each for 2,000 iterations. The training objective for base and layer Gaussian is the L1 loss and D-SSIM term between the ground-truth views and the rendered views. We use a single 80G A100 GPU for reconstruction and the reconstruction time for each layer is 1.5 minutes on average for $1024\times 1024$ resolution inputs.
4.2 Comparison Methods.
To evaluate the performance of our approach in the context of text-driven 3D panoramic scene generation, We compare with previous methods in two phases: 2D Panorama Generation and 3D Panoramic Scene Reconstruction. For 2D Panorama Generation, we choose four approaches to check the quality and creativity of 2D panorama: Text2light [15] adopts the VQGAN [53] structure to synthesize an HDR panorama image from text in a two-stage auto-regressive manner. Panfusion [11] designs a novel dual-branch diffusion model to generate a 360-degree image from a text prompt. Diffusion360 [43] directly uses the DreamBooth [54] training on the panorama dataset [55]. LDM3D-pano [12] extends LDM3D [56] to generate a realistic RGB panorama and its corresponding depth. For 3D Panoramic Scene Reconstruction, we select three methods to compare the 3D performance of scenes: LucidDreamer [4] starts from a single image and a text prompt to create a 3D-GS scene. Text2Room [6] generates a textured 3D mesh scene with preset camera trajectory. Dreamscene360 [16] creates panoramic 3DGS scene with full $360^{\circ}$ coverage from text prompts. Note that, since DreamScene360 does not open-source its official code, we reproduce its model for comparison.
4.3 Qualitative Comparison
2D Panorama Generation. We show some qualitative comparisons with several state-of-the-art panorama generation works in Figure 5. Text2light ignores style prompts due to being trained on a realistic HDRI dataset based on the VQGAN structure, and the components in the generated panorama are relatively simple. The results by PanFusion are ambiguous and low in quality, while it can adapt to style changes due to its operation in panorama and perspective domains. LDM3D-pano is insensitive to style prompts and the results are low-resolution. The instances generated by Diffusion360 appear to be of higher quality compared to the above, but do not have the inherent spherical structural property of panoramas, and the style seems to be the same. This is because the model is merely trained on the panorama dataset based on Dreambooth [54] without injecting any prior knowledge. Our method achieves the highest quality both in consistency and diversity, presenting creative and reasonable generations.
3D Panoramic Scene Reconstruction. To more precisely evaluate the quality of 3D panoramic scene reconstruction, we present qualitative comparisons with Text2Room [6], LucidDreamer [4], and DreamScene360 [16] across two dimensions. For full $360^{\circ}\times 180^{\circ}$ view consistency, we input an image with a text prompt to guide scene creation and render multiple views from the center of the scene. This approach synthesizes the texture of the panoramic environment, allowing for a more intuitive assessment after generation. As shown in Figure 6, LucidDreamer [4] and Text2room [6] fail to cover the full $360^{\circ}\times 180^{\circ}$ view, resulting in semantic incoherence and artifacts due to their successive inpainting-based strategy. DreamScene360 [16] supports a $360^{\circ}\times 180^{\circ}$ view at a single fixed viewpoint, but it deviates from the input conditions, and the quality of the generated results is relatively low. In contrast, our model excels in maintaining full $360^{\circ}\times 180^{\circ}$ view consistency while demonstrating superior content creativity. To evaluate the capability for free trajectory rendering, we design a zigzag trajectory to guide the camera’s movement through the scene, with novel view renderings sampled along the trajectory for comparison. As illustrated in Figure 7, we showed 3 random samples from this fixed flythrough trajectory. Compared with the other three methods, our model achieves a more complete 3D scene with consistent textures and a reasonable geometric structure.
Figure 6: Qualitative comparisons in full 360°×180° Scene. We compare the panorama and multiple views of the scene generated by four methods.  LayerPano3D  exhibits more consistent content with full $360^{\circ}\times 180^{\circ}$ coverage.
Figure 7: Qualitative comparisons in Scene Free Exploration. We show the novel view renderings along a zigzag trajectory to compare the capability of free trajectory rendering. Our method is able to maintain high-quality content rendering and does not show distortion or gaps in unseen space, which shows the ability of  LayerPano3D  to create hyper-immersive panoramic scenes.
Figure 8: Analysis on the Layer Completion Inpainting. We present the panorama inpainting results for three methods guided by the same text prompt: “empty scene, nothing” [47]. Our model effectively handles complex scenarios, delivering clear results with consistent and coherent structures.
4.4 Quantitative Comparison
2D Panorama Generation. We adopt three metrics for quantitative comparisons: 1) Intra-Style [57, 34] evaluates the coherence of the generated panoramas; 2) FID [58] evaluates both fidelity and diversity; 3) CLIP [59] measures the compatibility of results and input prompts. We selected 73 prompts to generate panoramas for each method and averaged the results for all metrics. Moreover, a user study is also conducted to further evaluate the quality of panoramas, where we project 4 views at a fixed FOV ($90^{\circ}$) to the user for sorting. As shown in Table I, our method achieves the best scores in both FID and CLIP metrics, indicating its high fidelity and strong compatibility with the input prompts. The Intra-Style results demonstrate that our model achieves global coherence across the image, consistently maintaining the overall style. While Text2light has a smaller Intra-Style score, this is primarily due to its tendency to generate monotonous results with extensive uniform color block backgrounds. We also attained outstanding results in aesthetic metrics, marginally trailing Diffusion360. However, as discussed in Section 4.3, panoramas generated by Diffusion360 lack the inherent spherical structural property, resulting in an unreasonable final perspective projection.
Moreover, we conducted a user study comparing generated panoramas, evaluating them across three dimensions: Coherence (internal consistency), Plausibility (plausibility of perspective views), and Compatibility (align with prompts). Following [60], we use the Average User Ranking (AUR) as a preference metric where users rank each result on a scale of 1 to 5 (lower is worse). We invite 29 users including graduate students that expertise in 3D and average users to rank the 60 results from 5 methods individually. The results in Table I demonstrate that our method outperforms others by a significant margin across all dimensions.
TABLE I: Quantitative comparison with SoTA methods on 2D Panorama Generation. Bold indicates the best result, and underline indicates the second-best result.
      Method       FID $\downarrow$       Intra-Style $\downarrow$       CLIP $\uparrow$       User Study (AUR)
      Coherence $\uparrow$       Plausibility $\uparrow$       Compatibility $\uparrow$
      Text2light [15]       286.90       0.31       18.69       2.95       2.47       1.89
      Panfusion [11]       283.80       18.66       21.22       3.06       3.03       3.26
      Diffusion360 [43]       274.03       3.70       21.65       3.08       3.22       3.24
      LDM3D-pano [12]       298.64       8.05       20.58       2.56       2.84       2.87
      Ours       255.60       2.24       22.08       3.35       3.44       3.73
TABLE II: Qualitative comparison with SoTA methods on 3D Panoramic Scene. Bold indicates the best result.
Method Appearance Geometry User Study (AUR)
NIQE $\downarrow$ BRISQUE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Roll-Mean $\downarrow$ Roll-Var $\downarrow$
$360^{\circ}\times 180^{\circ}$uparrow
Free-path $\uparrow$
Text2room [6] 5.117 44.598 30.437 0.890 0.033 1.877 1.566 1.76 2.28
LucidDreamer [4] 5.877 54.230 31.020 0.973 0.029 2.775 2.167 1.85 1.29
DreamScene360 [16] 5.083 39.403 28.925 0.941 0.050 1.696 2.524 2.94 2.90
Ours 4.374 39.006 40.797 0.980 0.012 0.540 0.028 3.45 3.53
3D Panoramic Scene Reconstruction. Following [2], we adopt two widely-used non-reference image quality assessment metrics, NIQE [61] and BRISQUE [62], to evaluate the capability of free trajectory rendering, i.e. the novel view quality along an immersive navigation path in the 3D scene. Furthermore, we also measure the 3D reconstruction quality on training views via classic reconstruction metrics: PSNR, SSIM and LPIPS [63]. In addition, we render four orthogonal views, each distributed at a fixed FOV ($90^{\circ}$), elevation ($0^{\circ}$) and four equidistant azimuths ($0^{\circ},90^{\circ},180^{\circ},270^{\circ}$) respectively, to predict perspective field. Building upon  [64], we independently predict camera parameters and evaluate the horizon tilt problem by calculating the Roll-Mean and Roll-Var, i.e. the mean and variance of the elevation angles across these views. As shown in Table II, our method surpasses the existing methods in both novel view quality metrics (NIQE and BRISQUE) and 3D reconstruction metrics (PSNR, SSIM, and LPIPS). Furthermore, we conducted another user study to further evaluate the quality of the generated 3D panoramic scene from two aspects: 1) $360^{\circ}\times 180^{\circ}$ view consistency and 2) free path rendering quality. For the first aspect, we render 60 frames in a 360-degree view at the 0-degree and 45-degree elevation respectively for evaluation. For the second aspect, we select the same trajectory as in Figure 7 to render navigation videos for evaluation. We invite 43 users including graduate students that expertise in 3D and average users to rank the 40 results from 4 methods individually. The average ranking is shown in Table II. Our LayerPano3D achieves the best performance in both $360^{\circ}\times 180^{\circ}$ view consistency and free path rendering quality among all four approaches.
Figure 9: Ablation on the Circular Blending Strategy. The circular blending strategy ensures the continuity between the leftmost and rightmost sides of the panorama, which preserves the crucial characteristic of 360-degree panoramas.
Figure 10: Ablation on the Gaussian Selector. Our Gaussian Selector resolves the optimization conflicts between layers by re-activating the Gaussians from the previous layer that blocks new assets. With the Gaussian Selector, the merged Gaussians are optimized to faithfully reconstruct the ground-truth panorama views.
4.5 Analysis and Ablative Study
Ablation on Circular Blending. Figure 9 demonstrates that the circular blending strategy effectively maintains geometric continuity, contributing to the generation of a seamless 360-degree panorama. Without this design, the rightmost and leftmost sides of the panorama could form inconsistent patterns, resulting in a discontinuous panorama.
Analysis on Layer Completion Inpainting. We discuss the effectiveness of our fine-tuned panorama inpainter in layer completion. We compare the inpainting results among three approaches: LaMa [42], Stable Diffusion inpainting model and our proposed inpainter. As illustrated in Figure 8, LaMa produces inconsistent texture and blurry artifacts at large-scale inpainting. Stable Diffusion tends to produce distorted new elements due to the domain gap between perspective and panoramic images. In contrast, thanks to the panorama finetuning, our module delivers clean inpainting results with coherent and plausible structures in the masked regions.
Ablation on Gaussian Selector. Our Gaussian selector is proposed to select the part of Gaussians that appears in the front of newly added scene assets. By selecting these Gaussians and re-activating them in the optimization, the model achieves accurate appearance and geometry at the current layer. As shown in Figure 10, the leftmost column is the scene Gaussians at layer 0. When adding the building assets at the first layer, the sky Gaussians from the previous layer partially block the building assets (right column). After using the Gaussian selector to select and optimize the sky Gaussians, these Gaussians learn to either be translucent and pruned for low opacity or move to be a part of the building assets. Therefore in the middle column, we observe a consistent scene with no obvious blockage of the new building assets thanks to the Gaussian Selector.
Figure 11: Analysis on Panorama Rendering at Off-center Viewpoints. We show that LayerPano3D ’s robustness in rendering $360^{\circ}\times 180^{\circ}$ consistent panorama at various viewpoints within a large range of space in the generated scene. Compared with our single-layer panorama baseline, LayerPano3D renderings exhibit high-quality content without any holes or gaps from occlusion.
Analysis on Panorama Renderings at Off-center Viewpoints. In Figure 11, we demonstrate that LayerPano3D is robust to render consistent panorama images at various locations besides the original camera location in the center. We sample four camera locations on circular trajectories on the hemisphere centered at the origin and render 24 views at ($-45^{\circ},0^{\circ},45^{\circ}$) elevation to compose new panorama images. By evaluating panorama renderings at new viewpoints, we show that our generated panoramic scene is $360^{\circ}\times 180^{\circ}$ consistent and enclosed, robust to various viewpoints at any angle. Compared to the single-layered 3D panorama, our multi-layered 3D panorama exhibits no gaps or holes from the scene occlusion, demonstrating our capability for free 3D exploration in the generated scenes.
5 Conclusion
In this paper, we propose LayerPano3D, a novel framework that generates hyper-immersive panoramic scene from a single text prompt. Our key contributions are two-fold. First, we propose the text-guided anchor view synthesis pipeline to generate detailed and consistent reference panorama. Second, we pioneer the Layered 3D Panorama representation to show complex scene hierarchies at multiple depth layers, and lift it to Gaussians to enable free 3D exploration. Extensive experiments show the effectiveness of LayerPano3D in generating $360^{\circ}\times 180^{\circ}$ consistent panorama at various viewpoints and enabling immersive roaming in 3D space. We believe that LayerPano3D holds promise to advance high-quality, explorable 3D scene creation in both academia and industry.
Limitations and Future Works. LayerPano3D leverages good pre-trained prior to construct panoramic 3D scene, i.e., panoramic depth prior for 3D lifting. Therefore, the created scene might contain artifacts from inaccurate depth estimation. With advancements in more robust panorama depth estimation, we hope to create high-quality panoramic 3D scenes with finer asset geometry.
Figure 12: Additional results of  LayerPano3D on Diverse Generation. LayerPano3D generates various hyper-immersive Layered 3D Panorama that cover cityscape, landscape and indoor environments.

### v2

@@
 LayerPano3D: Layered 3D Panorama for Hyper-Immersive Scene Generation
-Abstract
+Abstract.
 3D immersive scene generation is a challenging yet critical task in computer vision and graphics.
-A desired virtual 3D scene should 1) exhibit omnidirectional view consistency, and 2) allow for free exploration in complex scene hierarchies.
+A desired virtual 3D scene should 1) exhibit omnidirectional view consistency, and 2) allow for large-range exploration in complex scene hierarchies.
 Existing methods either rely on successive scene expansion via inpainting or employ panorama representation to represent large FOV scene environments.
@@
 Our key insight is to decompose a reference 2D panorama into multiple layers at different depth levels, where each layer reveals the unseen space from the reference views via diffusion prior.
-LayerPano3D comprises multiple dedicated designs: 1) we introduce a novel text-guided anchor view synthesis pipeline for high-quality, consistent panorama generation.
+LayerPano3D comprises multiple dedicated designs: 1) We introduce a new panorama dataset Upright360 , comprising 9k high-quality and upright panorama images, and finetune the advanced Flux model on Upright360 for high-quality, upright and consistent panorama generation related tasks.
 2) We pioneer the Layered 3D Panorama as underlying representation to manage complex scene hierarchies and lift it into 3D Gaussians to splat detailed 360-degree omnidirectional scenes with unconstrained viewing paths.
@@
 We believe that LayerPano3D holds promise for advancing 3D panoramic scene creation with numerous applications.
-Index Terms:
-3D Scene Generation, Panorama Generation, Diffusion Models, Neural Rendering
-Figure 1: Overview of LayerPano3D.
+More examples please visit our project page: ys-imtech.github.io/projects/LayerPano3D/
+journalyear: 2025
+Figure 1.
+Overview of LayerPano3D.
 Guided by simple text prompts, LayerPano3D leverages multi-layered 3D panorama to create hyper-immersive panoramic scene with $360^{\circ}\times 180^{\circ}$ coverage, enabling free 3D exploration among complex scene hierarchies.
-1 Introduction
+1. Introduction
 The development of spatial computing, including virtual and mixed reality systems, greatly enhances user engagement across various applications, and drives demand for explorable, high-quality 3D environments.
-We contend that a desired virtual 3D scene should 1) exhibit high-quality and consistency in appearance and geometry across the full $360^{\circ}\times 180^{\circ}$ view; 2) allow for free exploration among complex scene hierarchies with clear parallax.
-In recent years, many approaches in 3D scene generation [1, 2, 3] were proposed to address these needs.
-One branch of works [4, 5, 6, 7, 8, 9] seeks to create extensive scenes by leveraging a “navigate-and-imagine” strategy, which successively applies novel-view rendering and outpaints unseen areas to expand the scene.
+We contend that a desired virtual 3D scene should 1) exhibit high-quality and consistency in appearance and geometry across the full $360^{\circ}\times 180^{\circ}$ view; 2) allow for exploration among complex scene hierarchies with clear parallax.
+In recent years, many approaches in 3D scene generation (Gao et al., 2024; Li et al., 2024; Zhang et al., 2023b) were proposed to address these needs.
+One branch of works (Chung et al., 2023; Yu et al., 2023; Höllein et al., 2023; Fridman et al., 2023; Ouyang et al., 2023) seeks to create extensive scenes by leveraging a “navigate-and-imagine” strategy, which successively applies novel-view rendering and outpaints unseen areas to expand the scene.
 However, this type of approaches suffer from the semantic drift issue: long sequential scene expansion easily produces incoherent results as the out-paint artifacts accumulate through iterations, hampering the global consistency and harmony of the generated scene.
-Another branch of methods [10, 11, 12, 13, 14, 15] employs Equirectangular Panorama to represent $360^{\circ}$, large field of view (FOV) environments in 2D.
+Another branch of methods (Tang et al., 2023b; Zhang et al., 2024; Wang et al., 2022, 2023b; Chen et al., 2022) employs Equirectangular Panorama to represent $360^{\circ}$, large field of view (FOV) environments in 2D.
 However, the absence of large-scale panoramic datasets hinders the capability of panorama generation systems, resulting in low-resolution images with simple structures and sparse assets.
-Moreover, 2D panorama [10, 11, 13] does not allow for free scene exploration.
-Even when lifted to a panoramic scene [16], the simple spherical structure fails to provide complex scene hierarchies with clear parallax, leading to occluded spaces that cause blurry renderings, ambiguity, and gaps in the generated 3D panorama.
-Some methods [17] typically use inpainting-based disocclusion strategy to fill in the unseen spaces, but they require specific, predefined rendering paths tailored for each scene, limiting the potential for free exploration.
-To this end, we present LayerPano3D, a novel framework that leverages Multi-Layered 3D Panorama for full-view consistent and free exploratory scene generation from text prompts.
+Moreover, 2D panorama (Tang et al., 2023b; Zhang et al., 2024; Wang et al., 2022) does not allow for flexible scene exploration.
+Even when lifted to a panoramic scene (Zhou et al., 2024b), the simple spherical structure fails to provide complex scene hierarchies with clear parallax, leading to occluded spaces that cause blurry renderings, ambiguity, and gaps in the generated 3D panorama.
+Some methods (Zhou et al., 2024a) typically use inpainting-based disocclusion strategy to fill in the unseen spaces, but they require specific, predefined rendering paths tailored for each scene, limiting the potential for flexible exploration.
+To this end, we present LayerPano3D, a novel framework that leverages Multi-Layered 3D Panorama for explorable, full-view consistent scene generation from text prompts.
 The main idea is to create a Layered 3D Panorama by first generating a reference panorama and treating it as a multi-layered composition, where each layer depicts scene content at a specific depth level.
@@
 Our contributions are two-fold.
-First, to generate high-quality and coherent $360^{\circ}\times 180^{\circ}$ panoramas, we propose a novel text-guided anchor view synthesis pipeline.
-By finetuning a T2I model [18] to generate 4 orthogonal perspective views as anchors, we prevent semantic drifts during panorama generation, while ensuring a consistent horizon level across all views.
-Furthermore, the anchor views enrich the panorama by incorporating complex structures and detailed features derived from large-scale, pre-trained perspective image generators.
-Second, we introduce the Layered 3D Panorama representation as a general solution to handle occlusion for different types of scenes with complex scene hierarchies, and lift it to 3D Gaussians [19] to enable free 3D exploration.
+First, to generate high-quality and coherent $360^{\circ}\times 180^{\circ}$ panoramas, we curate a new dataset, namely Upright360 , consisting of 9k high-quality, upright panorama images, and finetune the advanced Flux (Labs, 2023) model with panorama LoRA on it for panorama generation and inpainting.
+This feed-forward pipeline prevents semantic drifts during panorama generation, while ensuring a consistent horizon level across all views.
+Second, we introduce the Layered 3D Panorama representation as a general solution to handle occlusion for different types of scenes with complex scene hierarchies, and lift it to 3D Gaussians (Kerbl et al., 2023) to enable large-range 3D exploration.
 By leveraging pre-trained panoptic segmentation prior and K-Means clustering, we streamline an automatic layer construction pipeline to decompose the reference panorama into different depth layers.
-The unseen space at each layer is synthesized with a finetuned panorama inpainter [11].
+The unseen space at each layer is synthesized with the Flux-based inpainting pipeline.
 Extensive experiments demonstrate the effectiveness of LayerPano3D in generating hyper-immersive layered panoramic scene from a single text prompt.
 LayerPano3D surpasses state-of-the-art methods in creating coherent, plausible, text-aligned 2D panorama and full-view consistent, explorable 3D panoramic environments.
-Furthermore, our framework streamlines an automatic pipeline without any scene-specific navigation paths, providing more user-friendly interface for non-experts.
+Furthermore, our framework does not require any scene-specific navigation paths, providing more user-friendly interface for non-experts.
 We believe that LayerPano3D effectively enhances the accessibility of full-view, explorable AIGC 3D environments for real-world applications.
-2 Related Works
-In recent years, advancements in AI, particularly deep learning, along with increased computational power and large datasets, have driven rapid development in 2D and 3D generation  [20].
-These technologies have revolutionized digital content creation in fields like entertainment, design, and virtual reality.
-2.1 3D Representation
-Early neural rendering methods [21, 22, 23, 24] typically incorporate a multi-layer perceptron (MLP) to model 3D scenes as continuous functions.
-NeRF [25], being the representative MLP-based method, achieves state-of-the-art novel view synthesis performance due to the volume rendering and inductive bias nature of MLPs.
-However, as the MLP needs to evaluate on large amount of points along every camera ray, the rendering process becomes extremely exhaustive.
-Subsequent methods [26, 27, 28, 29, 19] improve on faster training and rendering by employing sparse, advanced scene representation.
-Among these, 3D Gaussian Splatting (3DGS) [19] has recently emerged as a rising star with transformative real-time rendering and high-definition resolutions.
-It explicitly parameterizes the 3D scene with 3D Gaussians, where each Gaussian is optimized to represent a volume and projected to 2D for rasterization.
-The discrete nature of 3D Gaussians enables flexible scene editing and composition for 3DGS.
-Therefore, our approach employs 3DGS as scene representation to facilitate multi-layer scene creation.
-2.2 3D Scene Generation
-Due to the recent success of diffusion models, 3D scene generation has also achieved some development.
-Scenescape [7] and DiffDreamer [30], for example, explore perpetual view generation through the incremental construction of 3D scenes.
+2. Related Works
+2.1. 3D Scene Generation
+Due to the recent success of diffusion models (Tang et al., 2023a; Poole et al., 2022), 3D scene generation has also achieved some development.
+Scenescape (Fridman et al., 2023) and DiffDreamer (Cai et al., 2023), for example, explore perpetual view generation through the incremental construction of 3D scenes.
 One major branch of work employ step-by-step inpainting from pre-defined trajectories.
-Text2Room [6] creates room-scale 3D scenes based on text prompt, utilizing textured 3D meshes for scene representation.
-Similarly, LucidDreamer [4] and WonderJourney [5] can generate domain-free 3D Gaussian splatting scenes from iterative inpainting.
+Text2Room (Höllein et al., 2023) creates room-scale 3D scenes based on text prompt, utilizing textured 3D meshes for scene representation.
+Similarly, LucidDreamer (Chung et al., 2023) and WonderJourney (Yu et al., 2023) can generate domain-free 3D Gaussian splatting scenes from iterative inpainting.
 However, this line of work often suffer from the semantic drift issue, resulting in unrealistic scene from artifact accumulation and inconsistent semantics.
-While some other approaches [31, 3, 32] endeavor to integrate objects with environments, they yield relatively low quality of comprehensive scene generation.
-Recently, our concurrent works, DreamScene360 [16] and HoloDreamer [17] also employ panorama as prior to construct panoramic scenes.
+While some other approaches (Cohen-Bar et al., 2023; Zhang et al., 2023b; Vilesov et al., 2023) endeavor to integrate objects with environments, they yield relatively low quality of comprehensive scene generation.
+Recently, our concurrent works, DreamScene360 (Zhou et al., 2024b) and HoloDreamer (Zhou et al., 2024a) also employ panorama as prior to construct panoramic scenes.
 However, they only achieve the $360^{\circ}\times 180^{\circ}$ field of view at a fixed viewpoint based on a single panorama of low-quality and simple structure, and do not support free roaming within the scene.
-In contrast, our framework leverages Multi-Layered 3D Panorama representation to construct high-quality, fully enclosed scenes that enable unconstrained navigation paths in 3D scene.
-2.3 Panorama Generation
+In contrast, our framework leverages Multi-Layered 3D Panorama representation to construct high-quality, fully enclosed scenes that enable larger-range exploration paths in 3D scene.
+2.2. Panorama Generation
 Panorama generation methods are often based on GANs or diffusion models.
 Early in this field, with the different forms of deep generative neural networks, GAN-based panorama generation methods explore many paths to improve quality and diversity.
-Among them, Text2Light [15] focuses on HDR panoramic images by employing a text-conditioned global sampler alongside a structure-aware local sampler.
+Among them, Text2Light (Chen et al., 2022) focuses on HDR panoramic images by employing a text-conditioned global sampler alongside a structure-aware local sampler.
 However, training GANs is challenging and they encounter the issue of mode collapse.
 Recently, some studies have utilized diffusion models to generate panoramas.
-MVDiffusion [10] generates eight perspective views with multi-branch UNet but the resulting closed-loop panorama only captures the $360^{\circ}\times 90^{\circ}$ FOV.
-The image generated from MultiDiffusion [33] and Syncdiffusion [34] is more like a long-range image with wide horizontal angle as they do not integrate camera projection models.
-PanoDiff [35] can generate $360^{\circ}$ panorama from one or more unregistered Narrow Field-of-View (NFoV) images with pose estimation and controlling partial FOV LDM, while the quality and diversity of results are limited by the scarcity of panoramic image training data like most other methods [14, 36, 37].
+MVDiffusion (Tang et al., 2023b) generates eight perspective views with multi-branch UNet but the resulting closed-loop panorama only captures the $360^{\circ}\times 90^{\circ}$ FOV.
+The image generated from MultiDiffusion (Bar-Tal et al., 2023) and Syncdiffusion (Lee et al., 2023) is more like a long-range image with wide horizontal angle as they do not integrate camera projection models.
+PanoDiff (Wang et al., 2023a) can generate $360^{\circ}$ panorama from one or more unregistered Narrow Field-of-View (NFoV) images with pose estimation and controlling partial FOV LDM, while the quality and diversity of results are limited by the scarcity of panoramic image training data like most other methods (Wang et al., 2023b; Li and Bansal, 2023; Wu et al., 2024).
 In contrast, our model can generate Multi-Layered 3D Panorama for immersive, high-quality, and coherent scene generation from text prompts.
-3 Method
-Figure 2: Pipeline Overview of LayerPano3D.
-Our framework consists of three stages, namely reference panorama generation, multi-layer panorama construction and panoramic 3D scene optimization.
+3. Method
+Figure 2.
+Pipeline Overview of LayerPano3D.
+Our framework consists of two stages, namely multi-layer panorama construction and panoramic 3D scene optimization.
 LayerPano3D streamlines an automatic generation pipeline without any manual efforts to design scene-specific navigation paths for expansion or completion.
-The goal of our work is to create a panoramic scene guided by text prompts, that encompasses a complete $360^{\circ}\times 180^{\circ}$ field of view from various viewpoints within an extensive range in the scene, while allowing for unconstrained trajectory for immersive exploration.
-LayerPano3D consists three stages.
+The goal of our work is to create a panoramic scene guided by text prompts.
+This generated scene encompasses a complete $360^{\circ}\times 180^{\circ}$ field of view from various viewpoints within an extensive range in the scene, while allowing for immersive exploration along complex trajectories.
+LayerPano3D consists two stages.
 In Stage I (Sec.
-3.1), we propose a text-guided anchor view synthesis pipeline paired with panoramic outpainting to generate high-quality, consistent panorama as reference.
+3.1), we first generate the reference panorama from text prompt by finetuning flux (Labs, 2023) with panorama LoRA in our Upright360 dataset.
+With the reference panorama, we construct our Layered 3D Panorama representation by iterative layer decomposition, completion and alignment process.
 In Stage II (Sec.
-3.2), with the reference panorama, we construct our Layered 3D Panorama representation by iterative layer decomposition, completion and alignment process.
-In Stage III (Sec.
-3.3), the Layered 3D Panorama is lifted to 3D Gaussians in a cascaded manner to enable free 3D exploration.
-3.1 Reference Panorama Generation
-Text-Guided Anchor View Synthesis.
-We start by generating four orthogonal anchor views to establish the fundamental geometric structure and visual appearance.
-This is achieved by fine-tuning a pre-trained perspective T2I diffusion model.
-For data preparation, we construct a high-quality RGB panorama dataset consisting of 911 outdoor images sourced from the web and 9684 images from Matterport3D [38].
-For each panorama, we use llava-v1.5-7b [39] to generate captions as text prompts and project four orthogonal views without overlap at a fixed FOV ($60^{\circ}$), elevation ($0^{\circ}$) and four azimuths ($0^{\circ},90^{\circ},180^{\circ},270^{\circ}$).
-This setting can effectively avoid perspective distortion and fisheye effect, and allow further refinement of consistency in the subsequent synthesis stages.
-Additionally, we arrange these four views into a $2\times 2$ grid image in a fixed order for the ease of training and draw a white cross-line with a certain pixel size between the grids as a split.
-Compared with object-level multi-view generation, it is harder for scene-level generation to diffuse more views with high quality and consistency in a single pass.
-Multi-view cross-attention-based methods [10, 11] learn high-level semantics at the expense of overfitting to monotonous patterns due to the scarcity of panorama data.
-Other approaches [1] require huge computation and time in both training and inference stages and generate relatively simple scenes with limited assets.
-Therefore, inspired by [40], we directly fine-tune Stable-Diffusion-XL [18] to efficiently generate high-quality anchor views as image grids for subsequent panorama generation, free of complex regularization techniques.
-This text-guided anchor view synthesis pipeline can scale positively with the potency of the underlying T2I model and unlock the capability of the base model without additional knowledge through lightweight fine-tuning.
-Furthermore, unlike the clean background in the object-level generation, scene backgrounds are often diverse and complex.
-Thus, instead of simply initializing each sub-grid with a 2D Gaussian blob as is commonly done in object-based multi-view approaches, we initialize the image grids with random colored noise and draw a white cross-line between the grids during the inference process, mimicking the training data.
-Equirectangular Panorama Synthesis.
-Equirectangular Projection (ERP) is a method that maps a 3D sphere onto a 2D map.
-In ERP, the lines of longitude are mapped to vertical lines with constant spacing, and the lines of latitude are mapped to horizontal lines with constant spacing.
-This mapping relationship is neither area-preserving nor angle-preserving and the mapping of latitude leads to strong distortions in the polar regions.
-To obtain a detailed, consistent and plausible panorama, we first generate a $360^{\circ}\times 90^{\circ}$ FOV panorama to obtain the principle structure and content, then successively extend it to $360^{\circ}\times 180^{\circ}$ FOV panorama.
-With the text-guided anchor view synthesis, we have the anchor perspective images $[I_{anc}^{i}]_{i=1}^{4}\in\mathbb{R}^{h\times w\times 3}$ of a scene captured by four orthogonal camera in the horizontal plane.
-Based on the mapping rule and $60^{\circ}$ FOV, these anchor views can be directly projected to form the incomplete $360^{\circ}\times 60^{\circ}$ FOV panorama without overlap.
-Here, we assume that the elevation angle of the anchors are $\phi_{anc}=0^{\circ}$ with azimuth angles $\{\theta_{anc}^{i}\}=\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$.
-Then, we set a viewpoint sequence with $\{(\theta_{j},\phi_{j})\}_{j=1}^{n}$ that allows for a 360-degree rotation around the scene in $90^{\circ}$ FOV to project the perspective views $\hat{I_{j}}$ and masks $M_{j}$:
-$[\hat{I_{j}},M_{j}]=\mathcal{ERP}^{-1}(Pano,[\theta_{j},\phi_{j},FOV]),$ (1)
-where the $\mathcal{ERP}^{-1}$ is the inverse ERP from panorama to perspective views in 3D.
-We further apply a text-guided inpainter $\mathcal{F}_{inpaint}$ [41] to synthesize the missing pixels in the masked regions $M_{j}$ of the incomplete views $\hat{I_{j}}$.
-Note that, $\hat{I_{j}}$ is additionally pre-processed with LaMa [42] before inpainting to clean up the masked pixel distribution with contextual knowledge:
-$I_{j}=\mathcal{F}_{inpaint}(LaMa(\hat{I_{j}},M_{j})).$ (2)
-Through iterative projection and inpainting, the $Pano_{360\times 60}$ is extended to $Pano_{360\times 90}$.
-The extension from $Pano_{360\times 90}$ to $Pano_{360\times 180}$ is more difficult, as the polar regions exhibit most distorted patterns that cannot simply be synthesized through perspective image inpainting.
-Therefore, we fine-tune [43] on panoramic image-caption pairs for polar outpainting.
-Another challenge is to enforce continuity between the leftmost and rightmost sides of the synthesized image during outpainting, as it is a crucial characteristic [44] of 360 panoramas.
-Following  [14, 33], we adopt a circular blending strategy in both denoising and VAE decoding stage, as illustrated in Figure 3.
-A circular closed loop is created by extending the rightmost side of the panorama with the pixels in the leftmost parts.
-Per-pixel adaptive weights $W_{k}$ are utilized to blend the denoised patches mapped from extensions and from the leftmost parts to generate the leftmost panorama.
-Similarly, in the VAE tiled decoding stage, to avoid tiling artifacts, we ensure that each tile overlaps and blends together to form a smooth output due to each tile using a different decoder.
-Therefore, each pixel of the panorama is computed as a weighted average of all its mapped patch updates $Z_{k,t}$ at each denoising step.
-$\Psi(\mathcal{P}_{t})=\sum_{k}\frac{\mathcal{F}_{map,k}^{-1}(W_{k})}{\sum_{m}\mathcal{F}_{map,m}^{-1}(W_{m})}\odot\mathcal{F}_{map,k}^{-1}(\Phi(Z_{k,t})),$ (3)
-where $\mathcal{F}_{map,k}^{-1}$ is image space mapping from patch $Z_{k}$ to its corresponding latent in $\mathcal{P}$; $\Psi$ and $\Phi$ denote the diffusion process at panorama-level and patch-level.
-Finally, we can get a high-quality result as reference for the subsequent layered panorama construction.
-Figure 3: The circular blending Strategy.
-During the inference process, to synthesize a $360^{\circ}$ seamless panorama, adaptive weights are utilized to blend the leftmost part of the latent with the rightmost part in denoising and VAE decoding stage.
-3.2 Multi-Layer Panorama Construction
+3.2), the Layered 3D Panorama is lifted to 3D Gaussians in a cascaded manner to enable large-range 3D exploration.
+3.1. Multi-Layer Panorama Generation
 We introduce the Layered 3D Panorama representation based on the following assumption: “an enclosed 3D scene contains a background and various assets positioned in front of it”.
 In this regard, using Layered 3D Panorama for 3D scene generation is a general approach to handle occlusion for various types of scene.
-To create a complete scene from the reference panorama, we propose to decompose it into $N+1$ layers along the depth dimension.
-As shown in Figure 2, these layers, arranged from farthest to nearest, represent both the scene background (layer 0) and the layouts situated behind the observation point.
-Layer Decomposition.
-As shown in Figure 2, the reference panorama is decomposed by first identifying the scene assets and then cluster these assets in different layers according to depth.
-First, we employ an off-the-shelf panoptic segmentation model [45] pretrained on ADE20K [46] to automatically find all scene assets visible in the reference panorama.
+To create a complete scene, we first generate a high-quality reference panorama from a single text prompt and decompose it into $N+1$ layers along the depth dimension.
+As shown in Fig.
+2, these layers, arranged from the farthest to the nearest, represent both the scene background (layer 0) and the layouts behind the observation point.
+3.1.1. Reference Panorama Generation and Upright360 .
+A good panorama captures rich scene details, creating an immersive and comprehensive view.
+This depth of detail fosters a stronger connection and deeper comprehension of the scene.
+Unlike the advancement in standard image generation, panorama generation still faces quality gaps.
+To ensure upright and geometrically consistent scenes, we curate a new dataset, namely Upright360 , comprising high quality, upright panorama images, and finetune the Flux (Labs, 2023) with LoRA on the Upright360 dataset.
+This lightweight training approach achieves optimal performance even with limited high-quality panorama data and can be directly extended to subsequent tasks, for example panorama inpainting for layer completion.
+For data curation, we first collect around 15k raw panorama images: 9684 panorama image samples are collected from Matterport3D (Chang et al., 2017), 1824 images from the web, and 3592 synthetic panoramas generated from Blockadelabs.
+Based on this, we leverage GeoCalib (Veicht et al., 2024), a state-of-the-art single-perspective-image calibration method, to filter upright panoramas.
+Specifically, we employ the Equirectangular Projection (ERP), a mapping technique that projects a 3D sphere onto a 2D plane, to generate four perspective views from each panorama.
+These views are extracted at a fixed field of view (FOV) of 90°, an elevation of 0°, and four distinct azimuths (0°, 90°, 180°, 270°).
+We then use GeoCalib to calibrate the four views, computing their pitch and roll variances for filtering.
+Panoramas with variances exceeding 1.0 are classified as non-upright and excluded: $Var({pitch_{1:4}})>1.0,Var({roll_{1:4}})>1.0$, resulting in the creation of the final Upright360 dataset.
+This dataset comprises 9423 high-quality panoramas, rigorously filtered from the original collection.
+Building upon the Upright360 dataset, we fine-tune Flux to develop a panorama LoRA (Hu et al., 2021b) for reference panorama generation.
+3.1.2. Layer Decomposition.
+As shown in Fig.
+2, the reference panorama is decomposed by first identifying the scene assets and then cluster these assets in different layers according to depth.
+First, we employ an off-the-shelf panoptic segmentation model (Jain et al., 2023) pretrained on ADE20K (Zhou et al., 2017) to automatically find all scene assets visible in the reference panorama.
 A good layer decomposition requires that the layer assets share a similar depth level within layers and are distant from assets in other layers.
@@
 According to the depth values, the assets are clustered into $N$ groups from layer 00 to $N-1$ and are merged into layer masks to guide the subsequent layer completion.
-Layer Completion.
+3.1.3. Layer Completion.
 With the layer mask, we focus on completing the unseen content caused by asset occlusion.
-In order to synthesize background pixels instead of creating new elements, we finetune PanFusion [11] with the ERP-perspective aware cross-attention as the panoramic background inpainter.
-Specifically, at each layer, our model takes the layer mask $M_{l}$, the reference panorama, and the “empty scene, nothing” [47] prompt as input, and output coherent content at the masked area.
+In order to synthesize background pixels instead of creating new elements, we directly utilize the above trained panorama lora and integrate it into the Flux-Fill model to accomplish domain transfer, thereby employing it as a panoramic canvas inpainter.
+Specifically, at each layer, our model takes the layer mask $M_{l}$, the reference panorama, and the “empty scene, nothing” (Zhang and Agrawala, 2024) prompt as input, and output coherent content at the masked area.
 The inpainted panorama at layer $l$ is denoted $P_{l}$ and is used as supervision to the subsequent panoramic 3D Gaussian scene optimization.
-Note that, we additionally apply SAM [48] to extend the layer mask, based on the inpainted panorama from the previous layer, to eliminate unwanted new generations from inpainting.
-Moreover, to enable free rendering in 3D, where observers can examine scenes from varying distances, the unprocessed textures of distant assets may appear blurred as the observer approaches.
+Note that, we additionally apply SAM (Kirillov et al., 2023) to extend the layer mask, based on the inpainted panorama from the previous layer, to eliminate unwanted new generations from inpainting.
+Moreover, to enable large-range rendering in 3D, where observers can examine scenes from varying distances, the unprocessed textures of distant assets may appear blurred as the observer approaches.
 Therefore, distant layers require higher resolution to preserve texture details at different viewpoints.
-To address this, Super Resolution (SR) module [49] is employed to enhance the resolution of the layered panorama from layer 00 (background layer) to layer $N$ (reference panorama), achieving a 2×2$\times$2 × upscale in resolution.
+To address this, Super Resolution (SR) module (Yang et al., 2023) is employed to enhance the resolution of the layered panorama from layer 00 (background layer) to layer $N$ (reference panorama), achieving a 2×2$\times$2 × upscale in resolution.
 SR processing significantly improves the texture quality of distant objects, maintaining their visual clarity and texture details even when observed from a closer perspective.
-Layer Alignment.
+3.1.4. Layer Alignment.
 Given the Layered 3D RGB Panorama $[P_{l}]_{l=0}^{N}$, we perform the depth prediction and alignment to ensure consistency in a shared space.
-To begin with, we apply the 360MonoDepth [50], to first estimate the layer $N$ (reference panorama) as the reference depth $P_{depth}^{N}$.
-The predicted depth $P_{depth}$ inherently contains extremely small values from normalization, causing peak-like artifacts under the reference camera position.
-Consequently, we post-process the directly output disparity map $\hat{P}_{disp}$ with the following transformation to mitigate the geometry distortion:
-$D_{\text{bias}}$ $=\left(\max(\hat{P}_{\text{disp}})-\min(\hat{P}_{\text{disp}})\right)\ /\ 4,$ (4)
-$\hat{P}_{\text{depth}}$ $=\max(\hat{P}_{\text{disp}})-\hat{P}_{\text{disp}}+D_{\text{bias}},$
-$P_{\text{depth}}$ $=\hat{P}_{\text{depth}}\ /\ \max(\hat{P}_{\text{depth}}).$
-To align the layer depth in 3D space, we find it infeasible to simply compute a global shift and scale as in  [4, 6] due to the nonlinear nature of ERP.
-Therefore, we leverage depth inpainting model $\mathcal{F}_{depth}$ from [51] to directly restore depth values based on the inpainted RGB pixels.
+To begin with, we apply the 360MonoDepth (Rey-Area et al., 2022), to first estimate the layer $N$ (reference panorama) as the reference depth $P_{depth}^{N}$.
+Then, to align the layer depth in 3D space, we find it infeasible to simply compute a global shift and scale as in  (Chung et al., 2023; Höllein et al., 2023) due to the nonlinear nature of ERP.
+Therefore, we leverage depth inpainting model $\mathcal{F}_{depth}$ from (Liu et al., 2024) to directly restore depth values based on the inpainted RGB pixels.
 $\mathcal{F}_{depth}$ harnesses strong generalizability from large-scale diffusion prior and synthesizes inpainted depth values at an aligned scale with the base depth.
 We start from reference panoramic depth $P_{depth}^{N}$ to implement step-by-step restoration from layer $N-1$ to layer 00:
-$P_{depth}^{l}=\mathcal{F}_{depth}(P_{l},M_{l}\odot P_{depth}^{l+1}),$ (5)
+(1) $P_{depth}^{l}=\mathcal{F}_{depth}(P_{l},M_{l}\odot P_{depth}^{l+1}),$
 where, in layer $l$, the inpainted panorama $P_{l}$ and masked depth map $M_{l}\odot P_{depth}^{l+1}$ are provided as inputs to $\mathcal{F}_{depth}$ for restoration.
-3.3 Panoramic 3D Gaussian Scene Optimization
-Figure 4: Illustration of the Gaussian Selector.
-Given the new asset point cloud, the Gaussian Selector identifies the active Gaussians for next layer’s optimization.
-(a) Optimized Gaussians at layer $l$ partially blocks new asset point cloud in layer $l+1$.
-(b) By computing the viewing directions and the distance to camera, Gaussian Selector identifies the blocking Gaussians.
-(c) The marked Gaussians are labeled as active, and merged with the new points for optimization.
-(d) Showed a special case where the optimized Gaussian center is behind the new points, but also identifies as an active Gaussian because its volume scale overlaps with the points (case 1).
-3D Scene Initialization.
-To enable free 3D exploration, we lift the Layered 3D Panorama to 3D Gaussians [19], where the Gaussians are initialized from the layered 3D panoramic point clouds.
+3.2. Panoramic 3D Gaussian Scene Optimization
+Figure 3. Illustration of the Gaussian Selector. Given the new asset point cloud, the Gaussian Selector identifies the active Gaussians for next layer’s optimization.
+3.2.1. 3D Scene Initialization.
+To enable large-range 3D exploration, we lift the Layered 3D Panorama to 3D Gaussians (Kerbl et al., 2023), where the Gaussians are initialized from the layered 3D panoramic point clouds.
 Considering the intrinsic spherical structure of panorama, we can easily transform an equirectangular image $P\in\mathbb{R}^{H\times W\times 3}$ into 3D point cloud $S(\theta,\phi,P_{depth})$.
@@
 Then, the corresponding 3D coordinates $(X,Y,Z)$ from the depth value $P_{depth}(\theta_{u},\phi_{v})$ are derived as follows:
-$X$ $=P_{depth}(\theta_{u},\phi_{v})\cos{\phi_{v}}\cos{\theta_{u}},$ (6)
+(2) $X$ $=P_{depth}(\theta_{u},\phi_{v})\cos{\phi_{v}}\cos{\theta_{u}},$
 $Y$ $=P_{depth}(\theta_{u},\phi_{v})\sin{\phi_{v}},$
@@
 To speed up the calculation, we map the points into 3D grids, then remove all points within grids that have less than $\beta_{2}$ number of neighbors.
-3D Scene Refinement.
+3.2.2. 3D Scene Refinement.
 During scene refinement, we devise two types of Gaussian training schemes for varying scene content: the base Gaussian for reconstructing the scene background and the layer Gaussian for optimizing scene layouts.
@@
 We observe that the quality of the optimized scene is easily hampered by unaligned layers, and sometimes $\mathcal{F}_{depth}$ does fail to produce perfectly aligned layer depths.
-Gaussians at layer $l$ could span into unwanted depth levels and block assets in the subsequent layer, as illustrated in Figure 4(a).
+Gaussians at layer $l$ could span into unwanted depth levels and block assets in the subsequent layer, as illustrated in Fig.
+3(a).
 To handle this issue, we introduce the Gaussian selector module to detect these conflicted Gaussians, re-activate them from frozen, and optimize them away from the blockage.
-First, the selector computes the distance vector from the camera center $\mathbf{o}=(0,0,0)$ to each new point $\mathbf{p}$, as in Figure 4(b).
-The absolute distance from asset points $\mathbf{p}$ and scene Gaussians $\mathbf{g}$ to the camera is denoted as $d_{\mathbf{p}}$ and $d_{\mathbf{g}}$ respectively:
-$d_{\mathbf{p}}=||\mathbf{p}-\mathbf{o}||_{2},\quad d_{\mathbf{g}}=||\mathbf{g}-\mathbf{o}||_{2}.$ (7)
-By examining all Gaussians that on the same ray with asset points but at a closer distance: $\mathbf{p}\ /\ d_{\mathbf{p}}=\mathbf{g}\ /\ d_{\mathbf{g}},\quad d_{\mathbf{g}}<d_{\mathbf{p}},$ we mark them as active (Figure 4(c)).
-We also observe that despite the Gaussian center being behind the asset points, if the scaling of the Gaussian covers the point, then the Gaussian is still affecting the overall rendering quality.
-Hence, we update the gaussian distance as
-$d^{\prime}_{\mathbf{g}}=d^{\prime}_{\mathbf{g}}-\max\{\mathbf{s}(\mathbf{g})\},$ (8)
-where $\mathbf{s}$ is the scaling of the Gaussian $\mathbf{g}$.
-If $d_{\mathbf{g_{1}}}>d_{\mathbf{p_{1}}}$ and $d^{\prime}_{\mathbf{g_{1}}}<d_{\mathbf{p_{1}}}$ as in Figure 4(d), the referred Gaussian is also marked active.
+First, the selector computes the distance vector from the camera center $\mathbf{o}=(0,0,0)$ to each new point $\mathbf{p}$, as in Fig.
+3(b).
+The absolute distance from asset points $\mathbf{p}$ and scene Gaussians $\mathbf{g}$ to the camera is denoted as $d_{\mathbf{p}}$ and $d_{\mathbf{g}}$ respectively:$d_{\mathbf{p}}=||\mathbf{p}-\mathbf{o}||_{2},\quad d_{\mathbf{g}}=||\mathbf{g}-\mathbf{o}||_{2}$.
+By examining all Gaussians that on the same ray with asset points but at a closer distance: $\mathbf{p}\ /\ d_{\mathbf{p}}=\mathbf{g}\ /\ d_{\mathbf{g}},\quad d_{\mathbf{g}}<d_{\mathbf{p}},$ we mark them as active ( Fig.
+3(c)).
 For efficient memory storage and fast look-up, we hash the distance vectors into a 3D grid.
 The mapping function from vector coordinates to grid indices writes: $f(\mathbf{p})=\text{ceil}(\beta_{3}\log(\textbf{p}+1))$.
-Figure 5: Qualitative comparisons in Panorama Generation.
-We compare 2D panorama generation methods across three dimensions: Creativity (red), Resolution (green), and Spherical Structure (blue).
-LayerPano3D produces high-resolution, creative, and consistent panoramas while accommodating multi-style generation requirements.
-4 Experiments
-4.1 Implementation Details
-In the anchor view synthesis stage, we train the model starting from SDXL-base-1.0 [18] with a batch size of 16 and learning rate of $10^{-5}$ for only 10K iterations on the 11k data pairs.
-The training is done using 4 NVIDIA A100 GPUs for 7 hours.
-For inference, we initialize four $512\times 512$ grids with random colored noise, subsequently superimposing a white cross-line with a pixel width of 10 to ensure that the final inference result can be accurately segmented into four grids of size $507\times 507$.
-In the reference panorama synthesis stage, we leverage the Stable Diffusion Inpainting pipeline [41] from huggingface [52] as the inpainter and repurposed Diffusion360 [43] as the outpainter to synthesize $360^{\circ}\times 180^{\circ}$ FOV panorama.
-In the layered panorama construction stage, we employ OneFormer [45] to obtain the panoptic segmentation map for the reference panorama.
+4. Experiments
+4.1. Implementation Details
+In the layered panorama construction stage, we train the panorama LoRA starting from Flux (Labs, 2023) with a batch size of 1 and learning rate of $10^{-5}$ for 100K iterations on our curated Upright360 dataset.
+The training is done using 6 NVIDIA A100 GPUs for 3 days.
+For Layer Decomposition, we employ OneFormer (Jain et al., 2023) to obtain the panoptic segmentation map for the reference panorama.
 Background categories are manually determined (i.e.
@@
 Generally, we cluster all asset masks into $N=3$ layers via KNN and merge all masks within each layer to form a unified layer mask.
-With the obtained layer mask, we combine our fine-tuned panorama inpainting model with LaMa [42] to achieve multi-layer completion and apply 360MonoDepth [50] to predict the reference panorama depth.
+With the obtained layer mask, we integrate the above trained panorama LoRA into FLux-Fill model and combine with LaMa (Suvorov et al., 2021) to achieve multi-layer completion and apply 360MonoDepth (Rey-Area et al., 2022) to predict the reference panorama depth.
 In the 3D panoramic scene optimization stage, we lift the panorama RGBD into 3D point clouds.
@@
 We use a single 80G A100 GPU for reconstruction and the reconstruction time for each layer is 1.5 minutes on average for $1024\times 1024$ resolution inputs.
-4.2 Comparison Methods.
-To evaluate the performance of our approach in the context of text-driven 3D panoramic scene generation, We compare with previous methods in two phases: 2D Panorama Generation and 3D Panoramic Scene Reconstruction.
-For 2D Panorama Generation, we choose four approaches to check the quality and creativity of 2D panorama: Text2light [15] adopts the VQGAN [53] structure to synthesize an HDR panorama image from text in a two-stage auto-regressive manner.
-Panfusion [11] designs a novel dual-branch diffusion model to generate a 360-degree image from a text prompt.
-Diffusion360 [43] directly uses the DreamBooth [54] training on the panorama dataset [55].
-LDM3D-pano [12] extends LDM3D [56] to generate a realistic RGB panorama and its corresponding depth.
-For 3D Panoramic Scene Reconstruction, we select three methods to compare the 3D performance of scenes: LucidDreamer [4] starts from a single image and a text prompt to create a 3D-GS scene.
-Text2Room [6] generates a textured 3D mesh scene with preset camera trajectory.
-Dreamscene360 [16] creates panoramic 3DGS scene with full $360^{\circ}$ coverage from text prompts.
-Note that, since DreamScene360 does not open-source its official code, we reproduce its model for comparison.
-4.3 Qualitative Comparison
-2D Panorama Generation.
-We show some qualitative comparisons with several state-of-the-art panorama generation works in Figure 5.
-Text2light ignores style prompts due to being trained on a realistic HDRI dataset based on the VQGAN structure, and the components in the generated panorama are relatively simple.
-The results by PanFusion are ambiguous and low in quality, while it can adapt to style changes due to its operation in panorama and perspective domains.
-LDM3D-pano is insensitive to style prompts and the results are low-resolution.
-The instances generated by Diffusion360 appear to be of higher quality compared to the above, but do not have the inherent spherical structural property of panoramas, and the style seems to be the same.
-This is because the model is merely trained on the panorama dataset based on Dreambooth [54] without injecting any prior knowledge.
-Our method achieves the highest quality both in consistency and diversity, presenting creative and reasonable generations.
-3D Panoramic Scene Reconstruction.
-To more precisely evaluate the quality of 3D panoramic scene reconstruction, we present qualitative comparisons with Text2Room [6], LucidDreamer [4], and DreamScene360 [16] across two dimensions.
-For full $360^{\circ}\times 180^{\circ}$ view consistency, we input an image with a text prompt to guide scene creation and render multiple views from the center of the scene.
-This approach synthesizes the texture of the panoramic environment, allowing for a more intuitive assessment after generation.
-As shown in Figure 6, LucidDreamer [4] and Text2room [6] fail to cover the full $360^{\circ}\times 180^{\circ}$ view, resulting in semantic incoherence and artifacts due to their successive inpainting-based strategy.
-DreamScene360 [16] supports a $360^{\circ}\times 180^{\circ}$ view at a single fixed viewpoint, but it deviates from the input conditions, and the quality of the generated results is relatively low.
+4.2. Comparison Methods.
+To evaluate the performance of our approach in text-driven 3D panoramic scene generation domain, we compare with existing methods in two phases: 2D Panorama Generation and 3D Panoramic Scene Reconstruction.
+For 2D Panorama Generation, we compare the quality and creativity of 2D panorama with Text2light (Chen et al., 2022) (GAN-based HDR panorama generation), Diffusion360 (Feng et al., 2023) (diffusion-based text-to-panorama generation) and Panfusion (Zhang et al., 2024) (dual-branch diffusion-based generation).
+For 3D Panoramic Scene Reconstruction, we compare with Text2Room (Höllein et al., 2023) (iterative indoor scene expansion with textured mesh), LucidDreamer (Chung et al., 2023) (single-view scene generation with 3DGS), and Dreamscene360 (Zhou et al., 2024b) (text-guided panoramic 3DGS scene generation).
+4.3. Qualitative Comparison
+4.3.1. 2D Panorama Generation.
+We show some qualitative comparisons with several state-of-the-art panorama generation works in Fig.
+6.
+Text2Light (Chen et al., 2022) struggles to effectively interpret text prompt due to being trained on a realistic HDRI dataset based on the VQGAN structure, and the components in the generated panorama are relatively simple.
+The results by PanFusion (Zhang et al., 2024) are ambiguous and low in quality.
+While the instances generated by Diffusion360 (Feng et al., 2023) exhibit superior quality in comparison to the aforementioned methods, they lack intricate scene details and are prone to the generation of artifacts.
+In contrast, our method achieves the highest quality, presenting creative and reasonable generations.
+4.3.2. 3D Panoramic Scene Reconstruction.
+We present qualitative comparisons with Text2Room (Höllein et al., 2023), LucidDreamer (Chung et al., 2023), and DreamScene360 (Zhou et al., 2024b) across two dimensions.
+First, for full $360^{\circ}\times 180^{\circ}$ view consistency, we render multiple views from scene center point with single input image and text prompts.
+As shown in Fig.
+4, LucidDreamer (Chung et al., 2023) and Text2room (Höllein et al., 2023) fail to cover the full $360^{\circ}\times 180^{\circ}$ view, resulting in semantic incoherence and artifacts due to their successive inpainting-based strategy.
+DreamScene360 (Zhou et al., 2024b) supports a $360^{\circ}\times 180^{\circ}$ view at a single fixed viewpoint, but the quality of the generated results is relatively low.
 In contrast, our model excels in maintaining full $360^{\circ}\times 180^{\circ}$ view consistency while demonstrating superior content creativity.
-To evaluate the capability for free trajectory rendering, we design a zigzag trajectory to guide the camera’s movement through the scene, with novel view renderings sampled along the trajectory for comparison.
-As illustrated in Figure 7, we showed 3 random samples from this fixed flythrough trajectory.
-Compared with the other three methods, our model achieves a more complete 3D scene with consistent textures and a reasonable geometric structure.
-Figure 6: Qualitative comparisons in full 360°×180° Scene.
+Second, to evaluate novel path rendering, we design a zigzag trajectory to guide the camera’s movement through the scene, with novel view renderings sampled along the trajectory for comparison.
+Fig.
+5 shows 6 random samples from this fixed flythrough trajectory.
+Compared with all three methods, our model achieves a more complete 3D scene with consistent textures and a reasonable geometric structure.
+Figure 4.
+Qualitative comparisons in full 360°×180° Scene.
 We compare the panorama and multiple views of the scene generated by four methods.
-LayerPano3D  exhibits more consistent content with full $360^{\circ}\times 180^{\circ}$ coverage.
-Figure 7: Qualitative comparisons in Scene Free Exploration.
-We show the novel view renderings along a zigzag trajectory to compare the capability of free trajectory rendering.
+LayerPano3D  exhibits consistent and rich details across full $360^{\circ}\times 180^{\circ}$ coverage, while other methods show obvious inconsistencies and disorganized patterns in regions that deviate from the input view.
+Figure 5.
+Qualitative comparisons in Large-range Scene Exploration.
+We show the novel view renderings along a zigzag trajectory to compare the capability of large-range scene exploration.
 Our method is able to maintain high-quality content rendering and does not show distortion or gaps in unseen space, which shows the ability of  LayerPano3D  to create hyper-immersive panoramic scenes.
-Figure 8: Analysis on the Layer Completion Inpainting.
-We present the panorama inpainting results for three methods guided by the same text prompt: “empty scene, nothing” [47].
-Our model effectively handles complex scenarios, delivering clear results with consistent and coherent structures.
-4.4 Quantitative Comparison
-2D Panorama Generation.
-We adopt three metrics for quantitative comparisons: 1) Intra-Style [57, 34] evaluates the coherence of the generated panoramas; 2) FID [58] evaluates both fidelity and diversity; 3) CLIP [59] measures the compatibility of results and input prompts.
-We selected 73 prompts to generate panoramas for each method and averaged the results for all metrics.
-Moreover, a user study is also conducted to further evaluate the quality of panoramas, where we project 4 views at a fixed FOV ($90^{\circ}$) to the user for sorting.
-As shown in Table I, our method achieves the best scores in both FID and CLIP metrics, indicating its high fidelity and strong compatibility with the input prompts.
-The Intra-Style results demonstrate that our model achieves global coherence across the image, consistently maintaining the overall style.
-While Text2light has a smaller Intra-Style score, this is primarily due to its tendency to generate monotonous results with extensive uniform color block backgrounds.
-We also attained outstanding results in aesthetic metrics, marginally trailing Diffusion360.
-However, as discussed in Section 4.3, panoramas generated by Diffusion360 lack the inherent spherical structural property, resulting in an unreasonable final perspective projection.
-Moreover, we conducted a user study comparing generated panoramas, evaluating them across three dimensions: Coherence (internal consistency), Plausibility (plausibility of perspective views), and Compatibility (align with prompts).
-Following [60], we use the Average User Ranking (AUR) as a preference metric where users rank each result on a scale of 1 to 5 (lower is worse).
-We invite 29 users including graduate students that expertise in 3D and average users to rank the 60 results from 5 methods individually.
-The results in Table I demonstrate that our method outperforms others by a significant margin across all dimensions.
-TABLE I: Quantitative comparison with SoTA methods on 2D Panorama Generation. Bold indicates the best result, and underline indicates the second-best result.
-      Method       FID $\downarrow$       Intra-Style $\downarrow$       CLIP $\uparrow$       User Study (AUR)
-      Coherence $\uparrow$       Plausibility $\uparrow$       Compatibility $\uparrow$
-      Text2light [15]       286.90       0.31       18.69       2.95       2.47       1.89
-      Panfusion [11]       283.80       18.66       21.22       3.06       3.03       3.26
-      Diffusion360 [43]       274.03       3.70       21.65       3.08       3.22       3.24
-      LDM3D-pano [12]       298.64       8.05       20.58       2.56       2.84       2.87
-      Ours       255.60       2.24       22.08       3.35       3.44       3.73
-TABLE II: Qualitative comparison with SoTA methods on 3D Panoramic Scene. Bold indicates the best result.
+Table 1. Quantitative comparison with SoTA methods on 2D Panorama Generation. Bold indicates the best result.
+Method FID $\downarrow$ Aesthetic $\uparrow$ CLIP $\uparrow$ User Study (AUR) $\uparrow$
+Text2light 286.90 4.57 18.69 1.34
+Panfusion 283.80 4.78 21.22 2.38
+Diffusion360 274.03 5.07 21.65 2.52
+Ours 223.51 5.86 22.25 3.76
+Table 2.
+Qualitative comparison with SoTA methods on 3D Panoramic Scene.
+LayerPano3D  achieves high-quality reconstruction and novel view synthesis while maintaining upright panoramic scene compared to other methods.
 Method Appearance Geometry User Study (AUR)
-NIQE $\downarrow$ BRISQUE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Roll-Mean $\downarrow$ Roll-Var $\downarrow$
+NIQE $\downarrow$ BRISQUE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Pitch-Mean $\downarrow$ Pitch-Var $\downarrow$
 $360^{\circ}\times 180^{\circ}$uparrow
 Free-path $\uparrow$
-Text2room [6] 5.117 44.598 30.437 0.890 0.033 1.877 1.566 1.76 2.28
-LucidDreamer [4] 5.877 54.230 31.020 0.973 0.029 2.775 2.167 1.85 1.29
-DreamScene360 [16] 5.083 39.403 28.925 0.941 0.050 1.696 2.524 2.94 2.90
-Ours 4.374 39.006 40.797 0.980 0.012 0.540 0.028 3.45 3.53
-3D Panoramic Scene Reconstruction.
-Following [2], we adopt two widely-used non-reference image quality assessment metrics, NIQE [61] and BRISQUE [62], to evaluate the capability of free trajectory rendering, i.e.
-the novel view quality along an immersive navigation path in the 3D scene.
-Furthermore, we also measure the 3D reconstruction quality on training views via classic reconstruction metrics: PSNR, SSIM and LPIPS [63].
-In addition, we render four orthogonal views, each distributed at a fixed FOV ($90^{\circ}$), elevation ($0^{\circ}$) and four equidistant azimuths ($0^{\circ},90^{\circ},180^{\circ},270^{\circ}$) respectively, to predict perspective field.
-Building upon  [64], we independently predict camera parameters and evaluate the horizon tilt problem by calculating the Roll-Mean and Roll-Var, i.e.
-the mean and variance of the elevation angles across these views.
-As shown in Table II, our method surpasses the existing methods in both novel view quality metrics (NIQE and BRISQUE) and 3D reconstruction metrics (PSNR, SSIM, and LPIPS).
-Furthermore, we conducted another user study to further evaluate the quality of the generated 3D panoramic scene from two aspects: 1) $360^{\circ}\times 180^{\circ}$ view consistency and 2) free path rendering quality.
-For the first aspect, we render 60 frames in a 360-degree view at the 0-degree and 45-degree elevation respectively for evaluation.
-For the second aspect, we select the same trajectory as in Figure 7 to render navigation videos for evaluation.
-We invite 43 users including graduate students that expertise in 3D and average users to rank the 40 results from 4 methods individually.
-The average ranking is shown in Table II.
-Our LayerPano3D achieves the best performance in both $360^{\circ}\times 180^{\circ}$ view consistency and free path rendering quality among all four approaches.
-Figure 9: Ablation on the Circular Blending Strategy.
-The circular blending strategy ensures the continuity between the leftmost and rightmost sides of the panorama, which preserves the crucial characteristic of 360-degree panoramas.
-Figure 10: Ablation on the Gaussian Selector.
-Our Gaussian Selector resolves the optimization conflicts between layers by re-activating the Gaussians from the previous layer that blocks new assets.
-With the Gaussian Selector, the merged Gaussians are optimized to faithfully reconstruct the ground-truth panorama views.
-4.5 Analysis and Ablative Study
-Ablation on Circular Blending.
-Figure 9 demonstrates that the circular blending strategy effectively maintains geometric continuity, contributing to the generation of a seamless 360-degree panorama.
-Without this design, the rightmost and leftmost sides of the panorama could form inconsistent patterns, resulting in a discontinuous panorama.
-Analysis on Layer Completion Inpainting.
-We discuss the effectiveness of our fine-tuned panorama inpainter in layer completion.
-We compare the inpainting results among three approaches: LaMa [42], Stable Diffusion inpainting model and our proposed inpainter.
-As illustrated in Figure 8, LaMa produces inconsistent texture and blurry artifacts at large-scale inpainting.
-Stable Diffusion tends to produce distorted new elements due to the domain gap between perspective and panoramic images.
-In contrast, thanks to the panorama finetuning, our module delivers clean inpainting results with coherent and plausible structures in the masked regions.
-Ablation on Gaussian Selector.
+Text2room (Höllein et al., 2023) 5.231 46.127 30.126 0.882 0.038 2.029 1.724 1.69 2.31
+LucidDreamer (Chung et al., 2023) 5.822 52.102 36.108 0.954 0.026 2.813 2.189 1.81 1.31
+DreamScene360 (Zhou et al., 2024b) 5.051 39.891 30.056 0.958 0.062 1.328 2.018 2.86 2.77
+Ours 4.023 38.287 42.057 0.986 0.015 0.732 0.032 3.64 3.61
+4.4. Quantitative Comparison
+4.4.1. 2D Panorama Generation.
+We adopt three metrics for quantitative comparisons: 1) FID (Heusel et al., 2017) evaluates both fidelity and diversity; 2) Aesthetic (Schuhmann et al., 2022) evaluates the aesthetics of panorama; 3) CLIP (Hessel et al., 2021) measures the compatibility of results with input prompts.
+Moreover, a user study is also conducted to further evaluate the quality of panoramas, where we project 4 views at a fixed FOV ($90^{\circ}$) to the user for sorting.
+Here, we report the average Average User Ranking (AUR) (Zhang et al., 2023a), which is computed based on an integrated assessment of coherence, plausibility, aesthetics, and compatibility dimensions.
+As shown in Tab.
+1, our method achieves the best scores among all quantitative metrics and human evaluation, demonstrating its fidelity, alignment with text and overall consistency.
+4.4.2. 3D Panoramic Scene Reconstruction.
+Following (Zhou et al., 2024b), we adopt non-reference image quality assessment metrics, NIQE (Mittal et al., 2012b) and BRISQUE (Mittal et al., 2012a), to evaluate novel view quality along scene navigation paths.
+We also follow (Zhou et al., 2024b) to measure the rendering quality with PSNR, SSIM and LPIPS (Zhang et al., 2018).
+In terms of geometry evaluation, we render 4 orthogonal views ($90^{\circ}$ FOV, $0^{\circ}$ elevation and { $0^{\circ},90^{\circ},180^{\circ},270^{\circ}$ azimuths) and predict the Pitch-Mean and Pitch-Var (mean and variance of the elevation angles) with (Veicht et al., 2024) to evaluate whether the scenes are upright.
+As shown in Tab.
+2, our method surpasses the existing methods in both novel view quality metrics (NIQE and BRISQUE), 3D reconstruction metrics (PSNR, SSIM, LPIPS) and while ensuring the upright panoramic scene.
+Furthermore, we conduct another user study for 3D panoramic scene evaluation from two aspects: 1) $360^{\circ}\times 180^{\circ}$ view consistency and 2) novel path rendering quality.
+For the first aspect, we render 60 frames to cover 360-degree view at the 0-degree and 45-degree elevation respectively for evaluation.
+For the second aspect, we use the same trajectory as in Fig.
+5 to render navigation videos for evaluation.
+We invite 52 users including graduate students that expertise in 3D and average users to rank the 40 results from 4 methods.
+The average ranking is shown in Tab.
+2.
+Our LayerPano3D achieves the best performance in both $360^{\circ}\times 180^{\circ}$ view consistency and novel path rendering quality among all four approaches.
+Figure 6.
+Qualitative comparisons in Panorama Generation.
+LayerPano3D demonstrates superior capability in generating high-quality outputs with precise alignment to text prompt, outperforming other methods in fidelity and input adherence.
+4.5. Analysis and Ablative Study
+In this section and supp., we show the analysis and ablation on the Gaussian Selector (Sec.
+4.5.1), Multi-layer design (single vs.
+multi; Sec.
+4.5.2), Layer Gaussians representation (supp.), layer inpainting (supp.) and 3DGS optimization efficiency (supp.).
+4.5.1. Ablation on Gaussian Selector.
 Our Gaussian selector is proposed to select the part of Gaussians that appears in the front of newly added scene assets.
 By selecting these Gaussians and re-activating them in the optimization, the model achieves accurate appearance and geometry at the current layer.
-As shown in Figure 10, the leftmost column is the scene Gaussians at layer 0.
+As shown in Fig.
+7, the leftmost column is the scene Gaussians at layer 0.
 When adding the building assets at the first layer, the sky Gaussians from the previous layer partially block the building assets (right column).
@@
 Therefore in the middle column, we observe a consistent scene with no obvious blockage of the new building assets thanks to the Gaussian Selector.
-Figure 11: Analysis on Panorama Rendering at Off-center Viewpoints.
-We show that LayerPano3D ’s robustness in rendering $360^{\circ}\times 180^{\circ}$ consistent panorama at various viewpoints within a large range of space in the generated scene.
-Compared with our single-layer panorama baseline, LayerPano3D renderings exhibit high-quality content without any holes or gaps from occlusion.
-Analysis on Panorama Renderings at Off-center Viewpoints.
-In Figure 11, we demonstrate that LayerPano3D is robust to render consistent panorama images at various locations besides the original camera location in the center.
+Figure 7. Ablation on the Gaussian Selector. With the Gaussian Selector, the merged Gaussians are optimized to faithfully reconstruct the ground-truth panorama views.
+Figure 8.
+Analysis on Panorama Rendering at Off-center Viewpoints.
+Compared with the single-layer variant, LayerPano3D render $360^{\circ}\times 180^{\circ}$ consistent panorama at various off-center viewpoints without any holes or gaps from occlusion.
+4.5.2. Analysis on Panorama Renderings at Off-center Viewpoints.
+In Fig.
+8, we demonstrate that LayerPano3D is robust to render consistent panorama images at various locations besides the original camera location in the center.
 We sample four camera locations on circular trajectories on the hemisphere centered at the origin and render 24 views at ($-45^{\circ},0^{\circ},45^{\circ}$) elevation to compose new panorama images.
 By evaluating panorama renderings at new viewpoints, we show that our generated panoramic scene is $360^{\circ}\times 180^{\circ}$ consistent and enclosed, robust to various viewpoints at any angle.
-Compared to the single-layered 3D panorama, our multi-layered 3D panorama exhibits no gaps or holes from the scene occlusion, demonstrating our capability for free 3D exploration in the generated scenes.
-5 Conclusion
+Compared to the single-layered 3D panorama, our multi-layered 3D panorama exhibits no gaps or holes from the scene occlusion, demonstrating our capability for larger-range, complex 3D exploration in the generated scenes.
+5. Conclusion
 In this paper, we propose LayerPano3D, a novel framework that generates hyper-immersive panoramic scene from a single text prompt.
@@
 First, we propose the text-guided anchor view synthesis pipeline to generate detailed and consistent reference panorama.
-Second, we pioneer the Layered 3D Panorama representation to show complex scene hierarchies at multiple depth layers, and lift it to Gaussians to enable free 3D exploration.
+Second, we pioneer the Layered 3D Panorama representation to show complex scene hierarchies at multiple depth layers, and lift it to Gaussians to enable large-range 3D exploration.
 Extensive experiments show the effectiveness of LayerPano3D in generating $360^{\circ}\times 180^{\circ}$ consistent panorama at various viewpoints and enabling immersive roaming in 3D space.
@@
 With advancements in more robust panorama depth estimation, we hope to create high-quality panoramic 3D scenes with finer asset geometry.
-Figure 12: Additional results of  LayerPano3D on Diverse Generation. LayerPano3D generates various hyper-immersive Layered 3D Panorama that cover cityscape, landscape and indoor environments.